OpenStudy (anonymous):

Please help!! completing the square x^2-2x=3

OpenStudy (gorv):

x^2-2x=3 x^2-2*1*x=3

OpenStudy (gorv):

add 1 on both side

OpenStudy (gorv):

x^2-2x+1=3+1=4 (x-1)^2=4

OpenStudy (anonymous):

Thats it?